### v2

# Does Generative AI speak Nigerian-Pidgin?: Issues about Representativeness and Bias for Multilingualism in LLMs

Abstract
Nigeria is a multilingual country with 500+ languages. Naija is a Nigerian-Pidgin spoken by approx. 120M speakers in Nigeria and it is a mixed language (e.g., English, Portuguese, Yoruba, Hausa and Igbo). Although it has mainly been a spoken language until recently, there are now various platforms publishing exclusively in Naija such as Naija Wikipedia. However, it is hard to distinguish by non-native from a larger pidgin languages spoken across West Africa known as West African Pidgin English (WAPE)—which is more simplied and understandable by wider audience in Ghana, Nigeria, and Cameroon. BBC news platform publishes exclusively in WAPE to cater for several countries in West Africa. In our paper, we show through statistical analyses and Machine Translation experiments that these two creole varieties do not represent each other (i.e., there are linguistic differences in word order and vocabulary) and Generative AI operates only based on WAPE. In other words, Naija is under-represented in Generative AI, and it is hard to teach LLMs with few examples.
Does Generative AI speak Nigerian-Pidgin?:
Issues about Representativeness and Bias for Multilingualism in LLMs
David Ifeoluwa Adelani Mila - Quebec AI Institute McGill University, Canada CIFAR AI Chair david.adelani@mcgill.ca                        A. Seza Doğruözthanks: Equal contribution LT3, IDLab, Universiteit Gent, Gent, Belgium as.dogruoz@ugent.be
Iyanuoluwa Shode Bloomberg ishode@bloomberg.net                        Anuoluwapo Aremu Lelapa AI
1 Introduction
Between 16th-19th centuries, there were contacts between Europeans and non-Europeans outside Europe. In West Africa, contacts between English and West African languages leading to simplified and mixed languages with linguistic features from several languages. These new forms of languages were lingua francas (i.e., common or bridge languages) that served for a mutual understanding between speakers of different languages for various purposes (e.g., trade, plantation agriculture, mining) (Mufwene, 2024). "Pidgin" and "creole" are used to refer to these languages. Although there is a lack of agreement about the precise definitions and coverage of these terms, pidgin roughly refers to "speech-forms which do not have native speakers, and are therefore primarily used as a means of communication among people who do not share a common language" Muysken et al. (1995). On the other hand, creoles could be assumed as extended pidgins which are more established and have native speakers especially in urban environments Muysken et al. (1995).
Until recently, West African Pidgin English (WAPE) has been mainly a spoken language with many local varieties (e.g., Nigerian Pidgin, Ghanaian Pidgin English, Cameroonian Pidgin English).
Nigeria is a multilingual country in West Africa African hosting over 500 different languages spoken by approx. 220 million people across 371 ethnic tribes Eberhard et al. (2019). It is the sixth most populous country in the world and Africa’s most populous country (also referred to as “The Giant of Africa” because of its teeming population. English is the official language and acquired mostly through formal education in Nigeria Agbo and Plag (2020). The three major tribes in Nigeria with their respective languages include Hausa (spoken by 63M speakers), Igbo (27M speakers), and Yorùbá (42M speakers). Nigerian Pidgin (Naija) is a mix of English with local languages (e.g., Portuguese, Yorùbá, Igbo, Hausa) Balogun (2013a), Oyebola and Ugwuanyi (2023). Naija is widely spoken (approx. 120M speakers) as a first and second language Adelani (2022) around the Southern part of Nigeria (e.g., Lagos and Niger-Delta) with origins going back to the English-Creole Atlantic Krio language family. It is also adopted as the unifying and unofficial language for communication across ethnically diverse groups. According to some researchers (e.g., Muysken et al. (1995)), Naija has evolved into a creole over time and has now native speakers as well. However, it is still referred to as a pidgin among the locals.
Although Naija may have words that sound similar to English, their meanings may vary and there is no standardized orthography for Naija  (Marchal et al., 2021; Akande and Salami, 2021; Lin et al., 2024).
Despite the large number of speakers across West Africa countries, WAPE remained as a spoken language until 2017 when the British Broadcasting Company (BBC) launched a news website (West African Pidgin English). It aims to target the diversity of WAPE speakers across different countries.  [^1] https://www.bbc.com/news/world-africa-40975399 Since 2022, Naija is also accepted as one of the languages on Wikipedia  [^2] https://meta.wikimedia.org/wiki/Requests_for_new_languages/Wikipedia_Nigerian_Pidgin. Although they are mutually understandable, there are linguistic and social differences between the two written varieties. For example, WAPE on BBC resembles English in terms of word order and vocabulary (see example (1)) with a simplified grammar (i.e., lacking auxiliary "were"). However, Naija on Wikipedia has a different word order and vocabulary choice (e.g., ”moto” instead of a "car" and ”wund" instead of ”injured" or ”wounded").
Figure 1: WAPE Locations: West African countries whose official language is English (The Gambia, Sierra Leone, Liberia, Ghana and Nigeria), Cameroon (North-Western and South-Western Anglophone region), and Krios immigrants to Equatorial Guinea. Map obtained from Yakpo (2024).
Example (1)
  1. WAPE Two pesin in di car dey injured.
  2. Naija Na wund di two pesin get for di moto.
  3. English Translation: Two persons in the car were injured.
From the sociolinguistic perspective, WAPE is favored mostly by educated Nigerians (who also speak, read and write in English) whereas Naija is used mostly in everyday life and it is more accessible for larger groups regardless of their background (e.g., educational status) (Akande and Salami, 2012). In terms of availability on the Internet, there is much more data about WAPE than Naija (160K sentences vs. 25K sentences) (Ogueji et al., 2021).
There is a growing need for more research for multilingual and low resource languages Doğruöz and Sitaram (2022); Doğruöz et al. (2021) in Generative AI systems. This need is even enhanced for pidgin and/or creole languages due to their high numbers of speakers but lack of data Lent et al. (2022, 2023). However, there are also unresolved issues about to what extent the available data on the Internet represent the language spoken in real life Doğruöz et al. (2023). It is crucially important that the different genres of the same pidgin and creole languages are also represented in these systems to be inclusive and accessible for all speakers/users with diverse backgrounds (e.g., educated vs. less educated).
In our paper, we address these issues for WAPE and Naija in the Nigerian context with the following contributions. We introduce Warri  [^3] Warri is the name of a city in the South-South region of Nigeria popular for the use of Naija. as a new MT evaluation data set including written WAPE on BBC news website and written Naija on Wikipedia. We also had interviews with the Naija Wikipedia writers to understand their motives and regulations about written Naija.   [^4] We will release the dataset in CC-BY-4.0 license. Our paper is the first to systematically analyze the similarities and differences between WAPE and Naija. Through a Machine Translation experiment, we find that Generative AI models (e.g., GPT-4o and LLaMa 3.1 8B (Touvron et al., 2023)) are biased towards WAPE do not include Naija despite large numbers of speakers. Further analysis shows that LLMs are hard to teach with few examples (e.g., 5-shots) to generate in Naija.
2 Related Work and Links with Our Study
Available research on representativeness originates from corpus linguistics where it is important to include samples from different textual sources to have a balanced and representative (smaller) corpus reflecting the variation in the (larger) corpora Biber (1993). Similarly, Crowdy (1993) states the significance of representative sampling corpora to minimize bias and maximize the credibility and consistency of the linguistic analyses. Therefore, representative sampling encompasses a broad spectrum of language usage across various contexts, genres, and demographic factors.
Generative AI systems depend on the availability of large data sets on the Internet. However, this assumption does not consider the representativeness of the variation in the available data sets which is especially difficult for multilingual and low resource languages Doğruöz et al. (2023). While developing language technologies for multilingual and low resource languages, it is crucially important to be aware of the linguistic variation (e.g., WAPE and Naija) within these languages and aim for representing the variation in a balanced way to prevent potential bias.
To investigate potential bias in the Nigerian context, the first step is to establish to what extent WAPE and Naija are similar or different from each other linguistically.
3 Warri MT benchmark dataset for WAPE and Naija
Dataset Creole Domain Average length (Pidgin) TRAIN DEV TEST
Bible PCM religious 25.2 31,051 1,500 1,500
JW300 PCM religious 17.2 23,322 1,500 1,500
UD PCM spoken 10.6 6241 1,500 1,500
Mafand PCM news 25.0 4,790 1,484 1, 564
Warri (single-way) WAPE news (BBC) 20.8 5 - 500
Warri (multi-way) WAPE & PCM Wiki 21.3 5 - 500
Table 1: Warri and other datasets: Warri is only used for evaluation in zero or few-shot (e.g. 5) setting. Warri (multi-way) have the same sentences in both WAPE and Naija (PCM, the ISO 639-3 code) pidgins unlike Warri (single-way). We label each dataset based on the specified pidgin assigned by the creators of the dataset.
To establish the Warri data set, we used a portion of WAPE BBC news data, previously used in MasakhaNER dataset (Adelani et al., 2021b)—a named entity recognition dataset with available untokenized texts, and downloaded Naija Wikipedia data from the HuggingFace. [^5] We make use of the 20231101 version, https://huggingface.co/datasets/wikimedia/wikipedia After the data collection, we created the parallel data set in English by recruiting two bilingual speakers. They translated about $505$ sentences from WAPE BBC data and Naija Wikipedia data into English. In this way, we maintained high-quality datasets by preventing the translators from mixing the features of the two pidgins into one. However, this also introduced a new obstacle (i.e., comparison of two test sets from slightly different domains (news vs. Wikipedia).
To handle the domain related obstacle, we created a multi-way parallel dataset for Wikipedia domain. First, we asked a bilingual speaker to translate the Naija Wikipedia sentences into English. Then, we asked a professional translator (a different person), to translate the English sentences into WAPE following the BBC style of writing.
Table 1 provides the details of our new Warri dataset, containing a single-way parallel sentences translated by two native speakers from WAPE BBC to English, and the multi-way parallel sentences where the English sentences have parallel translation in both WAPE BBC and Naija Wikipedia. Our test set composed of $500$ sentences and the remaining five sentences were for few-shot/in-context learning for LLMs.
Other datasets in Naija
There are also other parallel translation datasets (i.e., Naija-English). We also perform an analysis and evaluation on them, and compare them to our Warri dataset. Table 2 provides one example each per dataset.
Lang. Sample Sentences (English & Pidgin)
Bible
English And the Word became flesh, and dwelt among us
Naija 1 Den di Word kon shange to pesin and e stay with us for dis world
Naija 2 Di Wọd kọm bikọm human bin an Im liv wit ọs
JW300
English What can we do to make wise use of our freedom?
Naija Wetin go help us use our freedom well?
MAFAND
English Each group is supposed to submit its needs
Naija Each group suppose bring di things wey dem need kom
UD
English And I love the job with all my heart
Naija And I love di job as in wit all my heart
BBC
English It is great - nothing is better than proving people wrong
WAPE E dey great - nothing better pass make you prove pipo wrong
Wikipedia
English He married one wife with 7 children.
Naija Na one wife im mari an dem don bon 7 pikin.
Table 2: Example of different styles of Pidgin used in different corpora
  1. (a)
    Bible: We found two Naija Bibles online, the first one was translated by Wycliffe Bible Translators, and is part of the freely available eBible corpus (Akerman et al., 2023), and Naija Wikipedia writers also assert that this conforms with Naija rather than WAPE. The other translation was created by the Mercy Christian Ministry International (MCMI),  [^6] https://nigerianpidgin-bible.yolasite.com/ which is written to be closely similar to an African languages in Nigeria including the use of underdot diacritics as they are used in the Igbo language (e.g., “Wọd” for “Word”). Table 2 shows an example of a sentence in the two writing styles which shows there is no standard in the style of writing Naija. We focus on our analysis on Wycliffe since the MCMI does not have the complete bible online. We divided the bible data into 31,051/1,500/1,500 TRAIN/DEV/TEST split.
  2. (b)
    JW300: Similar to the Bible, JW300 (Agić and Vulić, 2019) is based on religious texts, bible studies and missionary reports of Jehovah Witness ministry in various languages. JW300 covers 343 languages including Naija. We divided the data into 23,322/ 1,500/ 1,500 TRAIN/DEV/TEST split.
  3. (c)
    UD-Pidgin: This is based on universal dependecy (UD) project for Naija Caron et al. (2019). The data is based on transcript of audio dialog between two Naija speakers.  [^7] https://github.com/UniversalDependencies/UD_Naija-NSC. We divided the data into 6,241/ 1,500/ 1,500 TRAIN/DEV/TEST split.
  4. (d)
    MAFAND: This is based on the news domain. The news articles were obtained from English Daily Trust newspaper, and translated to Naija Adelani et al. (2022). We make use of the same split as the MAFAND corpus with 4,790/1,484/1,564. Unlike the other datasets, it can be considered as “general domain” similar to Wikipedia.
Aside from parallel corpora, large amount of WAPE unlabelled texts have been collected in literature  from BBC Ogueji et al. (2021) to train language models such as AfriBERTa. The AfriBERTa corpus more than 160,000 sentences. Other sources of data for Naija are often smaller e.g., Naija tweets Muhammad et al. (2022). However, in our analysis, we primarily focus on parallel data sources.
4 Experimental setup
single-way multi-way
Metric Bible JW300 UD MAFAND WAPE) Wiki (WAPE) Wiki (Naija)
Jaccard Similarity ([0,1] range)
Unigram $\uparrow$ 0.133 0.295 0.537 0.554 0.712 0.802 0.517
Bigram $\uparrow$ 0.025 0.086 0.149 0.178 0.289 0.371 0.167
Trigram $\uparrow$ 0.002 0.025 0.055 0.076 0.151 0.207 0.084
Levenshtein distance $\downarrow$ 88.5 56.5 30.3 58.0 26.6 21.6 53.6
BLEU $\uparrow$ 0.8 11.2 16.8 20.9 36.1 46.8 23.4
ChrF++ $\uparrow$ 20.5 30.4 43.7 54.7 65.2 73.4 51.3
BERTScore $\uparrow$ 72.4 79.6 82.6 82.4 87.4 90.5 79.8
Table 3: Lexical overlap and Levenshtein distance on Warri benchmark. Lexical overlap is measured by Jaccard similarity between English and WAPE (BBC) and Naija (Wikipedia). For multi-way Warri corpus, the source of the data is from Wiki. The WAPE translation is denoted as Wiki (WAPE) but the original text is in Naija.
We conduct three types of experiments to find out if WAPE and Naija888pcm is the ISO 639-3 code for Naija pidgins are similar to each other: (1) Statistical analysis of the texts obtained from different datasets to find out how similar they are to English and to each other. (2) Cross-corpus zero-shot transfer results when a machine translation (MT) model is trained on one dataset and evaluated on the another. We expect domains that are similar should have a higher performance Adelani et al. (2021a); Lee et al. (2022). Similarly, we expect transfer results to be higher if the creoles are similar in writing. (3) Prompting an LLM to find out whether WAPE or Naija is represented in Generative AI. We compare the result to the evaluation of Warri MT benchmark dataset when trained on Mafand.
4.1 Statistical analysis of the texts
First, we compute lexical similarity between the English portion of each dataset and Pidgin by measuring Jaccard similarity (in percentage) for each corpus unigram, bigram, and trigram tokens. Secondly, we compute the Levenshtein distance (Levenshtein, 1965) which is an edit distance between each English test sentences and their translations to WAPE and Naija pidgins. Finally, we make use of three additional text generation metrics to measure their similarity to English: BLEU (Papineni et al., 2002), ChrF++ (Popović, 2017) and BERTScore (Zhang et al., 2020). BLEU and ChrF++ are n-gram matching metrics, while BLEU focuses on word-level matching, ChrF++ helps with evaluating character-level differences which are more common for Pidgin, thus more reliable. BERTScore is an embedding-based metric that can measure the semantic relationship between the sentence embeddings of two sentences, thus, they provide better correlation with human judgements.
4.2 Cross-corpus zero-shot transfer results
Here, we evaluate the performance of training an MT model e.g., on MAFAND, and evaluate on Bible, JW300, UD, BBC, and Wikipedia test sets. We perform evaluation based on ChrF++ due to it’s reliability on capturing the character-level differences between Pidgin and English. Following Adelani et al. (2022), we leveraged a pre-trained model to train an MT model by fine-tuning M2M-100 (418M) on each source data, and evaluated on the remaining test sets of our datasets. In our experiment, only four datasets have training data: Bible, JW300, UD, and MAFAND. The Warri benchmark is only for evaluation.
4.3 Prompting of LLMs
We prompted GPT-4o  [^9] GPT-4o pre-training data is up to December 2023. and LLaMa 3.1 8B & 70B Dubey et al. (2024) to generate translations in either Pidgin or English in both zero-shot or few-shots settings (with one or five examples). A sample prompt is provided in Appendix A. The prompting result is compared with the supervised training of MT models on the Mafand dataset which is also in the general domain.
5 Experimental Results
5.1 Statistical analysis results
In Table 3, by computing a lexical similarity between the $n$-gram tokens of each genre, we show that WAPE in both news (BBC) and Wikipedia domains consistently have a higher Jaccard similarity score with its parallel English corpus for all $n$-grams, compared to other datasets with the Naija Pidgin label. For example, the unigram similarity score for WAPE was around $0.712-0.802$ while the others are much lower between $0.133$ (Bible) and $0.554$ (Mafand). Interestingly, we find UD, Mafand and Warri Wiki having similar Jaccard similarity.
Furthermore, Levenshtein distance provides an additional evidence of a clear difference between WAPE and Naija. We find that it takes more than twice edit-distance to transform the English sentences to Naija (Wiki) than to WAPE (BBC) and WAPE (Wiki). Naija (Wiki) requires more edits in characters, which shows that it is farther from English compared to the WAPE pidgin. In other words, these two pidgins are quite different than each other linguistically. Similarly, we find longer Levenshtein distance for other datasets: JW300 Mafand, and Bible with $56.5$, $58.0$ and $88.5$ respectively. On the otherhand, UD dataset has a shorter Levenshtein distance compared to others which we attribute to the shorter utterances of the dataset (see Table 1).
Finally, our experiments on text generation metrics such as BLEU, ChrF++ and BERTScore shows that WAPE (BBC) is more similar to English than any of the other Pidgin datasets we evaluated. We find BLEU to be less reliable for this evaluation, achieving only $0.8$ for the Bible while ChrF achieve relatively higher score. We attribute this to several character-level differences between the Bible Pidgin and the English. In general, we find higher score for both WAPE (BBC) and WAPE (Wiki) achieving $65.2-73.4$ ChrF++, while Naija (Wiki) achieved $51.3$. BERTScore evaluation also confirmed this reaching a score of $90.5$ for WAPE while Naija achieved $79.8$.
Figure 2: Cross-corpus transfer results: Evaluation based on ChrF++
Single-way (news) Multi-way parallel (Wiki)
WAPE (BBC) WAPE (Wiki) Naija (Wiki)
Evaluation Task BLEU ChrF++ BERTScore BLEU ChrF++ BERTScore BLEU ChrF++ BERTScore
wape/pcm $\rightarrow$ en
0-shot: Mafand $\rightarrow$ Warri 57.6 76.3 94.5 68.6 83.4 96.2 35.0 59.1 86.7
0-shot: LLaMa 3.1 8B 51.5 74.8 92.1 58.5 79.5 93.6 37.8 64.5 89.0
0-shot: LLaMa 3.1 70B 56.1 76.4 93.5 63.6 81.6 94.3 42.7 67.7 90.7
0-shot: GPT-4o 59.3 78.8 94.4 65.5 83.6 96.3 43.5 68.9 92.2
en $\rightarrow$ wape/pcm
0-shot: Mafand $\rightarrow$ Warri 54.7 75.2 91.6 61.0 79.5 92.9 26.5 51.8 83.6
0-shot: LLaMa 3.1 8B 41.3 66.8 88.4 45.0 68.5 89.9 22.9 48.0 81.9
1-shot: LLaMa 3.1 8B 41.0 67.9 87.9 44.0 72.1 87.7 26.1 50.2 82.3
5-shot: LLaMa 3.1 8B 49.2 72.0 90.2 53.4 76.4 91.2 26.4 50.7 83.2
0-shot: LLaMa 3.1 70B 46.1 69.8 89.7 43.8 67.8 89.5 25.4 50.9 83.2
1-shot: LLaMa 3.1 70B 50.6 73.1 90.8 56.0 76.9 92.0 29.0 53.4 84.2
5-shot: LLaMa 3.1 70B 58.1 77.2 92.1 61.5 80.3 93.2 28.0 53.1 84.7
0-shot: GPT-4o 51.8 72.3 91.4 53.8 74.8 92.0 26.7 51.7 83.1
1-shot: GPT-4o 58.7 76.9 92.6 57.7 79.7 92.8 29.6 54.3 84.6
5-shot: GPT-4o 63.5 79.6 93.2 64.9 83.1 93.8 30.0 54.7 85.1
Table 4: Evaluation on Warri dataset: single-way and multi-way parallel (same sentences translated to both genres) test sets: We compared the performance of MT to different genres using GPT-4-Turbo and adapted M2M-100 (418M) from Mafand training set.
5.2 Machine translation evaluation
While statistical analysis already provided the difference between WAPE and Naija, evaluation on a task provides a different perspective i.e., would a model trained on WAPE, perform well on Naija, and vice versa? Also, what is the transfer performance of a model trained on one creole to another especially for text generation tasks?
Figure 2 shows our results on MT. The transfer performance depends on both the similarity of the domains (e.g. religious vs. news) and similarity of the creole style of writing. For example, religious datasets such as Bible and JW300 generally transfers poorly to other domains. Similarly, we find Mafand and UD in the news and spoken conversation domains also do not transfer well to religious domains.
If we focus on both Mafand and UD datasets, we find that they often have a higher zero-shot transfer result to WAPE (BBC and Wiki) than to Naija (Wiki). In the English-Pidgin direction, Mafand achieved between $76-80$ ChrF++ on WAPE while achieving only $52$ ChrF++ on Naija (PCM). Surprisingly, we find Mafand transfering better to WAPE than to its own test set showing the simplicity of generating WAPE pidgin compared to Naija. We have similar observation when transfering from UD. Our evaluation in the Pidgin-English also confirm this hypothesis that translating from WAPE to English is easier for MT systems than Naija.
5.3 LLM performance on Warri Benchmark
Here, we focus on finding out which pidgin is represented in the current LLMs, and whether they support several pidgin variants which are accessible for different communities of Naija speakers. We evaluated the performance of LLM in translating from and into WAPE and Naija.
Mafand MT model and LLMs represent WAPE more
[…]
Method Translation
English Rimi said he gave them the money because they did all that was expected of them to do in their different areas.
Reference (Naija) Rimi sey im giv dem di moni bikos dem "folo do evritin wey dem sopos du for dia difren aria.
Reference (WAPE) Rimi tok say im give dem di money becos dem do evritin wey dem suppose do for dia different areas.
English-Naija translation
0-shot: GPT-4o Rimi talk say e give dem the money because dem do everything wey dem suppose do for their different areas.
1-shot: GPT-4o Rimi tok say im give dem di moni because dem do all di tin wey dem suppose do for dia different area.
5-shot: GPT-4o Rimi tok se im dash dem di moni becos dem do all di tins wey dem suppose do for dia diffren areas.
0-shot: LLaMa 3.1 70B Rimi talk say him give dem money because dem do all wetin dem suppose do for dia different area.
1-shot: LLaMa 3.1 70B Rimi tok say im give dem moni because dem do all wet dem suppose do for dia different areas.
5-shot: LLaMa 3.1 70B Rimi talk say im dash dem di moni bikos dem do wetin dem suppose do for dia different areas.
English-WAPE translation
0-shot: GPT-4o Rimi talk say e give dem di money because dem do all wey dem suppose do for their different areas.
1-shot: GPT-4o Rimi talk say e give dem the money because dem do all wey dem expect make dem do for their different areas.
5-shot: GPT-4o Rimi talk say e give dem di money because dem do all wey dem expect dem to do for their different areas.
0-shot: LLaMa 3.1 70B Rimi tok say him give dem money because dem do all wetin dem suppose do for dia different areas.
1-shot: LLaMa 3.1 70B Rimi talk say im give dem moni because dem do all wetin dem suppose do for dia different areas.
5-shot: LLaMa 3.1 70B Rimi talk say im give dem di money because dem do all wet dem expect dem to do for dem different areas.
Table 5: Qualitative analysis on Predicted translations on Warri dataset: multi-way parallel output. Words/phrases expressed in Naija are in violet color, WAPE words are in cyan, while English words that ought to be translated are in red.
Can we teach LLMs different genres with only a few examples?
Our result (Table 4) of prompting GPT-4o, LLaMa 3.1 8B and LLaMa 3.1 70B shows that providing one or five examples is effective in providing more boost in performance for generating Pidgin sentences. [^10] The exact 1-shot and 5-shots examples are provided in the Appendix D For GPT-4o, the performance improved over zero-shot result by $+4.9$ ChrF++ when the LLM is prompted with one example translation of WAPE, and $+8.3$ ChrF++ when prompted with five examples, on the multi-way test set. However, the boost in performance is very small when Naija (Wiki) examples are provided. The boost is only $+2.6$ and $+3.0$ when one example and five examples are provided during the prompting of GPT-4o. This clearly shows that GPT-4o is more biased toward the WAPE than Naija and it is difficult to teach LLM with few examples. We believe one of the reason for this performance difference is because the WAPE BBC is the largest unlabelled data available on the web (Ogueji et al., 2021). Other sources that are more representative of Naija, are often in smaller quantity like Wikipedia. We observe similar trend for the LlaMa models where LLaMa 3.1 70B attained up to $80.3$ ChrF++ with 5-shots ($-2.7$ points when compared to GPT-4o), while LLaMa 3.1 8B achieved $76.4$.
We provide a qualitative example in Table 5, where we show that with one or five examples, the LLaMa 3.1 70B LLM slightly changes its writing style to be more similar to Naija but sometimes the model combine the vocabulary of WAPE and Naija. For example, in the 5-shot translation of “…was expected of them to do in their different areas”, LLaMa 3.1 70B translated it to be “suppose do for dia different areas” which is more similar to the WAPE translation of the same sentence, while in Naija the words of “suppose”, “different” and “area” are spelt differently: “sopos du for dia difren aria”. On the other hand, GPT-4o almost produced an accurate translation in Naija except the use of “suppose” rather than “sopos” and “becos” (a WAPE word) instead of “bikos”. This implies that the model is able to learn in-context, however, it is still biased to WAPE without few-shot examples. With more examples, we may be able to teach the model Naija with supervised fine-tuning of instruction data containing Naija-English parallel sentences. Qualitative examples for WAPE, shows that the LLMs are able to correctly generate sentences in zero-shot setting without additionally few shot examples which confirms our hypothesis that the LLMs are biased to WAPE.
6 Qualitative interviews with Naija Wikipedia writers
To validate our study, we interviewed two native speakers of Naija who contribute to the writing and editing of Naija Wikipedia articles. Some Wikipedian contributors have a public profile with links to their email addresses and social media accounts like Twitter or Linkedln. We sent emails to two Naija contributors with public profile and conducted a one hour interview with each of them.  [^11] We provided honorarium of \$$\$11$ to each interviewee.
Through these interviews, we found that Naija Wikipedia writers are not linguists or language experts but they are just volunteers without a formal linguistic training. They have a passion for their Naija and make effort toward establishing a writing system which is very similar to the way they passive people speaking it in their community or in Nigeria, rather than trying to reach a wider audience like BBC. (2) They made a lot of effort to create standardized way of writing during the Naija Wikipedia incubator programme, otherwise, Wikipedia would not have included it as a separate language. These volunteer writers were part of the Wikipedia incubator program from the start, and they are part of the editors team of Naija Wikipedia. This team mentors new contributors about how to write the Naija reflecting the patterns how Naija is spoken (sometimes with a few adjustments to make it readable since original spoken Naija form could be different than English (e.g. “moto” instead of “car”)).
The writers also mentioned that they read the available literature (e.g., Ofulue and Esizimetor (2010)) about Naija and follow the recommended rules by Naija linguists  Balogun (2013b); Aghoghovwia et al. (2010) before starting to contribute to Naija Wikipedia.
In terms of content of the Naija articles, the writers focus on biographies of notable people (e.g., musicians and actors) in Nigeria. They are not allowed to contribute to sensitive topics like health except it is a direct translation from an high-resource language like English. The target audience for their articles is Naija speakers in Nigeria rather than the whole West Africa (targeted by BBC). To achieve this, they prefer words that come from local Nigerian languages (e.g., Hausa, Igbo and Yoruba), which many Nigerians are familiar with rather than words commonly understandable across West Africa.
7 Conclusion
Different versions of Pidgins mixed with English and local languages are used in West Africa but not all of them have standardized writing systems. Since 2017, BBC internet broadcasts in WAPE targets West African countries with the goal of reaching a wider audience in a standardized writing style.
Like many countries in West Africa, Nigeria is a multilingual country with both richness and challenges that come along with the diversity in the linguistic landscape. Although the official language is English, Naija is a lingua franca that brings speakers of different Nigerian languages together regardless of their linguistic, social or educational backgrounds. Since 2022, it is also a written language on Wikipedia with the efforts of volunteers.
Although both pidgin varieties are used in Nigeria, we prove that WAPE and Naija are very different from each other linguistically and current Generative AI models are built upon WAPE only. This is probably due to more availability of data on the Internet for the WAPE pidgin rather than Naija.
It is well-known that the lack of data on low resource languages is a key challenge for current AI systems which operate mostly on and for low resource languages. In our paper, we prove that the situation is much more challenging for linguistically rich areas (e.g., West Africa) where the pidgin varieties with the most data on the Internet gets represented on AI systems and the others are not visible. This could lead to bias towards favoring language preferences of certain speakers/users instead of being more inclusive toward the users/speakers of other pidgins.
8 Limitation
There are few limitations of our work (1) Our evaluation dataset is small, although we argue that 500 may be good enough as a test set for MT, however, we only have maximum of 5 sentences we could use for few-shot learning or in-context learning. Moreover, with additional sentences (e.g. 2.5K-5K parallel sentences as recommended in (Adelani et al., 2022)), we may be able to adapt M2M-100 model to produce better generation of the Wikipedia genre. (2) Our analysis is limited to one task which is machine translation, we hope to extend this analysis to other tasks in the future. (3) our analysis only cover Naija spoken in Nigeria, we hope to extend our analysis to other English-based creole in West Africa like Ghananian Pidgin, Cameronian Pidgin, and Krio in the future.
Appendix A Prompt Template
Appendix B MAFAND training
We fine-tune MAFAND dataset on M2M-100 (418M) using the same hyparameters stated in Adelani et al. (2022) i.e. number of training epochs of 10, batch size of 32, source and target maximum sequence length of 200, and beam size of 10.
Appendix C Licence of Warri
We plan to release it publicly under the CC-4.0-NC due to BBC portion of the dataset that cannot be for commercial use. However, Warri (multi-way) has a licence of CC-4.0 international.
Appendix D Five shot examples provided
Prompt
Task Description You are a helpful assistant who is an expert in translating English sentences to Pidgin using two varieties: West African Pidgin English and Nigerian-Pidgin, I would provide you with five examples of the different varieties, your task is to follow the style of the writing of the specified variety when translating the sentences.
Example Example 1:
English: Innocent Ujah Idibia was born on 18 September 1975, that is well known as 2baba, a Nigerian singer, songwriter, producer, philantropist.
West African Pidgin English: Innocent Ujah Idibia wey dem born for 18 September 1975, wey dem know as 2baba, be a Nigerian singer, songwriter, producer, philantropist.
Nigerian-Pidgin: Innocent Ujah Idibia (dem bon am for 18 September 1975), wey pipul no wel wel as 2baba, na Naija singa, songraita an podusa an im sabi dash pipul moni an gift wel wel.
Example Example 2:
English: He was born in Jos, Nigeria
West African Pidgin English: Dem born am for Jos, Nigeria
Nigerian-Pidgin: Dem bon am for Jos for inside Naija.
Example Example 3:
English: He is from the Idoma ethnic group
West African Pidgin English: Im be from di Idoma ethnic group
Naija: Im na Idoma pesin.
Example Example 4:
English: Idoma is in the southern part of Nigeria
West African Pidgin English: Na southern part of Nigeria Idoma dey
Nigerian-Pidgin: Idoma dey for di south side for Naija.
Example Example 5:
English: Before July 2014, he used 2face Idibia as his stage name
West African Pidgin English: Before July 2014, i dey use 2face Idibia as im stage name
Nigerian-Pidgin: Bifor July 2014 na 2face Idibia bi di nem wey im dey yuz for stej.
Prompt ’Translate this sentence to Nigerian Pidgin
Input Alexander Abolore Adegbola Akande was born on 17 January 1980, well known as 9ice, a Nigerian singer, dancer, and songwriter.
Output: Alexander Abolore Adegbola Akande (dem bon am for 17 January 1980), wey pipul sabi well well as 9ice, na Naija singa, dansa, an songraita.
Table 6: Prompt template used for MT. An example prediction by GPT-4o